You have the option to elect to receive your $100,000 bonus in either Jan 2020 or deferring it into Jan 2021, for exactly 1 year. Your tax rate in 2020 is 30% and is expected to be 20% in 2021. The rate of return from investing your 2020 bonus would be 6%. Which year would provide the larger after tax result? Provide calculations. What other issues might you want to consider?
Answer #1

After tax amount if received in 2020 = 100,000(1-30%) = $70,000

Amount in 2021 after investment = 70,000(1.06) = $74,200

If received in 2021 = 100,000(1-20%) = $80,000

Hence, 2021 would provide larger after tax result

Other issues are:

1.requirement of money

2. Certainty of tax rates